

********************************************************************
UARS WEEKLY STATUS REPORT - 27 OCT 2000
********************************************************************
***********
SOLSTICE
***********
This week SOLSTICE supported the low spacecraft power condition during the
yaw maneuver by powering off during orbit eclipse and taking data only on
the daylit side of each orbit. Normal operations resumed on day 299.
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is not affected by the larger than normal pointing errors
since they are within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
However, level2 processing is halted due to deficiencies in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem and plan to have
a patch released by today (Oct 27).
***********
SUSIM
***********
SUSIM turned off for the yaw-around on 10/21/00 and was turned back on
and resumed taking data on 10/25/00. Otherwise, it continues to gather
a normal set of daily solar scans scheduled for predicted periods of
UARS telemetry.
We are preparing to begin the reprocessing the SUSIM UV spectral
irradiances to data version V21 in the next few weeks.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE was powered beginning-of-day today. All engineering parameters are
within nominal operating region. Gain and bore sight verification tests
will be conducted over the beta cycle.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is complete to
3 December 2000. The MPG continues to coordinate with the Flight
Operations Team and to assist the Project Scientist with the
scheduling of instrument operations.
***********
CDHF
***********
HIGHLIGHTS
----------
o UOAS (Solstice requested) delivery should be received at the CDHF sometime
next week. We will notified all PI SW reps with release notes.
o UARS PCG continue to notify PI Software Representatives regarding Data Gaps
that occur in the LEVEL0 data due to the UARS Spacecraft Anomaly.
o DAAC DATA:
- MLS Version 5 data is being previewed and loaded into the DAAC Transfer
Tables. Updating of various fields (per 5 day profile) in the DAAC Tables
that are needed to meet DAAC criteria have been done. We are verifying.
Additional updating appears to be required to reflect north/south
variations in min /max lat values. DAAC (James Johnson) has been
notified when to load.
- WINDII requested Version 11 be sent to the DAAC from start of mission to
31-Aug-1997. WINDII needs to update the UARS_QUALITY field. Currently
waiting response from WINDII that the fields have been updated. DAAC
will be notified.
o UARS CDHF REHOST:
- SUSIM v7.2 software representative Ed Esfandiari notified parallel
testing completed. *SUSIM responded with concerns regarding the output
product which is being investigated.
- HRDI v4.4 software representative Jim Raines was notified parallel
testing completed. HRDI has responded.
- GPI V4.05 tested successfully against the HRDI V4.4 Level 3AL data.
Frank Huang was requested to validate the GPI output product. *Frank
responded with questions and some concerns.
- SOLSTICE v4.62 software representative PI Barry Knapp has successfully
validated the parallel testing output.
- HALOE v4.82 software representative Lance Deaver was notified the
parallel test days completed. *HALOE has responded that they see no major
differences in the output product, but have 3 more species to check out.
Currently waiting response.
- WINDII V5.11A software representative Brian Solheim was notified the
parallel test days are complete. WINDII will be requested to validate
the output product and compare against same days processed on current
system.
- PEM is ready to run.
***********
FOT
***********
REPORTING PERIOD
This report covers 14 October 2000 (Orbit 49697, GMT Day 288) through 20
October 2000 (Orbit 49801, GMT Day 294).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 290/2349 49741 Off
HALOE 292/2354 49771 Off
PEM 294/2317 49800 Off
The beta angle ranged between the angles of 34.7 degrees to 8.9 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
2 degrees on 23 October 2000 (DOY 297).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.6 to 1.0 degrees C. Additional battery performance data for each
battery on 13 October 2000 and 20 October 2000 are provided in Appendix D.
The clock error ranged from -5.2 to 8.3 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 19
¥ HALOE = 01
¥ WINDII = 01
¥ PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 05
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/221 - 2287.496937
· Transponder B: 00/244 - 2287.499127
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/293 show the observatory data loss to be 298 hours, 39
minutes, 55.664 seconds (an increase of 4 minutes 38.448 seconds since the last
report period). This is a 0.39 percent data loss which equals a 99.61 percent
data capture for the mission.
The increase of 4 minutes 38.448 seconds occurred over one (1) day:
· 4 minutes 38.448 seconds on Day 287 due to late acquisition caused by weather
conditions at White Sands.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Testing continues daily to get the TAM and FSS attitude control solution
operational. The UARS testing and training simulator (UTTS) is being used for
all testing. In order to insure the validity of the test the FOT has been
debugging some problems with the simulator attitude models. This week two bad
bits were found in the FSS data on the simulator. The FOT is attempting to
isolate the problem.
The plan for the 23 October 2000 yaw maneuver was made this week. It was
decided that for the maneuver the nominal script will be followed with a few
additions. The forward and reverse gyro biases will be uplinked before and
after the maneuver, respectively. This will be done to hold a proper attitude
without the FSS. The FSS will also be put into monitor mode and once the
residuals show reasonable values the FSS will be put back into update mode. We
are confident that this plan will be sufficient to maintain good attitude
control. If needed FDF will be standing by to provide support and contingency
procedures have been generated. All additions have been verified on the
simulator.
MEETINGS
The FOT attended a meeting to discuss UARS attitude control and yaw maneuver
plans on 16 and 18 October 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew has been rescheduled for October 23, 2000 (DOY 297).
This maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing). The maneuver was delayed so that maximum support from Flight
Dynamics, CDHF and the FOT would be present should an anomaly occur.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 20 Oct 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
49699 288/0449 00-568 NBTR NOOPS D-288
49707 288/1747 00-568 NBTR NOOPS D-288
49707 288/1829 00-569 NBTR NOOPS D-289
49714 289/0548 00-569 NBTR NOOPS D-289
49722 289/1750 00-569 NBTR NOOPS D-289
49722 289/1830 00-570 NBTR NOOPS D-290
49729 290/0551 00-570 NBTR NOOPS D-290
49736 290/1652 00-571 OBC GYRO BIAS
49737 290/1752 00-570 NBTR NOOPS D-290
49741 290/2349 00-573 WINDII Power OFF
49748 291/1112 00-572 NBTR NOOPS D-291
49750 291/1514 00-574 OBC GYRO BIAS
49752 291/1833 00-575 NBTR NOOPS D-292
49760 292/0636 00-575 NBTR NOOPS D-292
49765 292/1516 00-576 OBC GYRO BIAS
49770 292/2314 00-577 NBTR NOOPS D-293
49771 292/2354 00-578 HALOE Power OFF
49775 293/0639 00-577 NBTR NOOPS D-293
49780 293/1434 00-579 OBC GYRO BIAS
49782 293/1801 00-577 NBTR NOOPS D-293
49796 294/1436 00-580 OBC GYRO BIAS
49800 294/2317 00-581 PEM Power OFF, HEPS, ZEPS
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 20 Oct 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 20 Oct 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-None this report period
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 20 Oct 2000
13 Oct 2000 - GMT Day 287 Beta = 37.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +16.8/-67.2 +5.47/+4.51 +28.8/-15.2 25.0 1.030 21.0
3 +5.6/-140.0 +4.51/+3.23 +28.4/-15.2 25.0 1.030 20.9
20 Oct 2000 - GMT Day 294 Beta = 12.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +22.4/-5.6 +4.51/+3.87 +29.2/-12.8 25.8 1.040 19.4
3 +00.0/-95.2 +3.55/+2.60 +28.4/-12.8 25.8 1.040 19.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 20 Oct 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-044 FOT 49715 289/0735 23031 Dropout at WS
00-045 FOT 49719 289/1248 N/A TAC 6 Crashed
00-042 FOT 49773 293/0414 14170 Generic Late Acq.
00-042 FOT 49776 293/0749 14170 Generic Late Acq.
00-042 FOT 49800 294/2213 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-032 FHST 48211 188/2242 N/A FHST 2 Anomaly
00-040 FOT 48854 231/2111 22831 Data Dropout
00-044 FOT 49715 289/0735 23031 Dropout at WS
00-045 FOT 49719 289/1248 N/A TAC 6 Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-042 FOT see list Oct 00 14170 Generic Late Acq.
SEVNTFW00/42